OpenStudy (kaitlyn_nicole):

Let f(x) = 2x + 2. Solve f^−1(x) when x = 4. 1 3 4 10

OpenStudy (kaitlyn_nicole):

i need an explanation, i'm really bad at this lol

OpenStudy (anonymous):

this is exactly the same as being asked: "solve for \(x\) if \(2x+2=4\)"

OpenStudy (anonymous):

it is just worded funny is all

OpenStudy (kaitlyn_nicole):

thats the inverse?

OpenStudy (kaitlyn_nicole):

so x=1?

OpenStudy (anonymous):

no one but a math teacher writes "solve for \(f^{-1}(x)\) if \(x=4\) must more natural to say "find \(f^{-1}(4)\)

OpenStudy (anonymous):

yes

OpenStudy (kaitlyn_nicole):

oo ok, can you help me with another to make sure i get it?

OpenStudy (anonymous):

if \[f(x)=2x+2\] then \[f(1)=4\] so \[f^{-1}(4)=1\]

OpenStudy (anonymous):

sure

OpenStudy (kaitlyn_nicole):

Let f(x) = 2x − 6. Solve f−1(x) when x = 2. −5 4 −2 7

OpenStudy (kaitlyn_nicole):

ok so, it would be 2x-6=2

OpenStudy (anonymous):

yes

OpenStudy (kaitlyn_nicole):

and x=4

OpenStudy (anonymous):

yes

OpenStudy (kaitlyn_nicole):

thank you soooo much!

OpenStudy (anonymous):

yw, you did almost all the work!

OpenStudy (kaitlyn_nicole):

i have a couple more but ill post in another question

OpenStudy (anonymous):

k